OpenStudy (anonymous):

3. Find the exact value of tan ( 75 ).

OpenStudy (anonymous):

\(75=30+45\) use addition angle formula

OpenStudy (anonymous):

im confused...

OpenStudy (aravindg):

\[\tan(A+B)=\dfrac{\tan A+\tan B }{1-\tan A \times \tan B}\]

OpenStudy (aravindg):

here A=30 B=45

OpenStudy (aravindg):

can u do tht now? @JNaruto5796

OpenStudy (anonymous):

im working on it! thanks so much!!

OpenStudy (anonymous):

You could also use: \[ \tan(x+y) = \frac{\sin(x+y)}{\cos(x+y)} = \frac{\sin(x)\cos(y)+\sin(y)\cos(x)}{\cos(x)\cos(y)-\sin(x)\sin(y)} \]If you didn't remember the \(\tan(x+y)\) identity.

OpenStudy (anonymous):

wait do i put it in the calculator at all?

OpenStudy (aravindg):

you nevr need a caluclator for these !

OpenStudy (anonymous):

No, angles like 30 and 45 are not hard to find.

OpenStudy (aravindg):